# FedLPA: One-shot Federated Learning with Layer-Wise Posterior Aggregation

### v3

@@ 1 Introduction
  • We are the first to train global model parameters via constructing a multi-variate linear objective function and optimizing its quadratic form, which allows us to formulate and solve the problem in a convex form efficiently. Furthermore, from the theoretical analysis, we show that our proposed FedLPA has a linear convergence rate, ensuring good performance.
@@ 2.1 Federated learning on non-iid data
Previous work FedAvg [9] first introduced the concept of FL and presented the algorithm, which achieved competitive performance on i.i.d data, in comparison to several centralized techniques. However, it was observed in previous works [27, 28] that the convergence rate and ultimate accuracy of FedAvg on non-IID data distributions were significantly reduced, compared to the results observed with homogeneous data distributions.
@@ 2.2 One-shot federated learning
One-shot Federated Learning (FL) is an emerging and promising research direction characterized by its minimal communication cost. In the first study on one-shot FL [13], the approach involved on the aggregation of local models, forming an ensemble to construct the final global model. Subsequently, knowledge distillation using public data was applied in the following step. FedKT [14] brought forward the concept of consistent voting to fortify the ensemble. Recent research endeavors [19, 24] proposed data-free knowledge distillation schemes tailored for one-shot FL. These methods adopted the basic ensemble distillation framework as FedDF [26]. XorMixFL [22] introduced the use of exclusive OR operation (XOR) for encoding and decoding samples in data sharing. It is important to note that XorMixFL assumed the possession of labeled samples from a global class by all clients and the server, which might not align with practical real-world scenarios. A noteworthy innovation of DENSE [24] was its utilization of a generator to create synthetic datasets on the server side, circumventing the need for a public dataset in the distillation process. Co-Boosting [39] improves the ensemble when doing the distillation to improve the performance. FedOV [15] delved into addressing comprehensive label skew cases. FEDCVAE [23] confronted this challenge by transmitting all label distributions from clients to servers. These schemes [22, 14, 19, 24, 23, 15] exposed some client-side private information, leading to additional communication overhead and potential privacy leakage, e.g., FEDCVAE [23] needed all the client label distribution to be transmitted to the server side and FedOV [15] needed the clients to know the labels which were unknown. Instead, MA-Echo [40] adopted a unique approach by emphasizing the addition of norms among layer-wide parameters during the aggregation of local models. The project [41] focused on the theoretic analysis of the error in its approximation method. However, their method grappled with limited experiments and lacked detailed explanations of the approach. FedDISC [20], on the other hand, relied on the pre-trained model CLIP from OpenAI, where their reliance might not always align with practicality or suitability for diverse scenarios.
@@ 3.1 Objective formulation
Previous works [16, 17] give a common formula of the global posterior which consists of local posteriors $p(\bm{\theta}|\mathcal{D}_{k})$ under variational inference formulation. More details are discussed in Appendix B:
@@ 3.2 Approximating posteriors
where $\bm{\theta}^{*}$ is the optimal parameter vector, $\bar{\mathbf{H}}=\mathbb{E}_{s\in\mathcal{D}}[\mathbf{H}]$ is the average Hessian of the negative log posterior over a dataset $\mathcal{D}$. It is reasonable to approximate global and local posteriors as multi-variates Gaussian distributions with expectations $\bar{\bm{\mu}}=\bm{\theta}^{*}$ and $\bm{\mu_{k}}=\bm{\theta}_{k}^{*}$; co-variances $\bar{\mathbf{\Sigma}}=\bar{\mathbf{H}}^{-1}$ and $\mathbf{\Sigma}_{k}=\bar{\mathbf{H}}_{k}^{-1}$ [44]. The details are discussed in Appendix C.
@@
As a result, if given local expectation $\bm{\mu}_{k}$ and local co-variance $\mathbf{\Sigma}_{k}$, the global posterior is determined by Eq. 2 as below, the details are discussed in Appendix D.
@@
Therefore, the local co-variance $\mathbf{\Sigma}_{k}$ can be approximated by the empirical Fisher $\tilde{\mathbf{F}}_{k}$ [48, 49]. the details is discussed in Appendix E:
@@
The works [50, 17, 51] ignore co-relations between different parameters and only consider the self-relations of parameters as computing all co-relations is impossible. Thus, their methods are inaccurate. Detailed discussions and the novelty compared to previous works are in Appendix F.
In order to capture co-relations between different parameters efficiently, previous works [46, 43] estimate a block empirical Fisher information matrix $\mathbf{F}$ instead of assuming parameters are independent and approximating the co-variance by the diagonal of the empirical Fisher. As pointed out, co-relations inner a layer are much more significant than others [46, 52, 53], while computing the co-relations between different layers brings slight improvement but much more computation. Therefore, assuming parameters are layer-independent is a good trade-off. As a result, the approximated layer-wise empirical Fisher is block-diagonal, while the details are shown in Appendix G. For layer $l$ on client $k$, its empirical Fisher $\mathbf{F}_{k_{l}}$ is one of the diagonal blocks in the whole empirical Fisher for the local model and is factored into two small matrices as below,
@@
where $\otimes$ is the Kronecker product; $\mathbf{A}_{k_{l}}=\mathbb{E}\left[\hat{\mathbf{a}}_{k_{l-1}}\hat{\mathbf{a}}_{k_{l-1}}^{\top}\right]+\pi_{l}\sqrt{\lambda}\mathbf{I}$ and $\mathbf{B}_{k_{l}}=\mathbb{E}\left[\hat{\mathbf{b}}_{k_{l}}\hat{\mathbf{b}}_{k_{l}}^{\top}\right]+\frac{1}{\pi_{l}}\sqrt{\lambda}\mathbf{I}$ are two expectation factor matrices over the data samples; $\hat{\mathbf{a}}_{k_{l}}$ is the activations and $\hat{\mathbf{b}}_{k_{l}}$ is the gradient of the pre-activations of layer $l$ on client $k$, $\lambda$ is the hyperparameter and $\pi_{l}$ is a factor minimizing approximation error in $\mathbf{F}_{k_{l}}$ [46, 49, 54]. $\mathbf{A}_{k_{l}}$ and $\mathbf{B}_{k_{l}}$ are symmetric positive definite matrices [45, 46]. The details are discussed in Appendix G.
@@ 3.4 Estimating the global expectation
As shown in Eq. 11, obtaining the global expectation $\bar{\bm{\mu}}_{l}$ requires calculating the inverse of $\bar{\mathbf{\Sigma}}_{l}^{-1}$ as Eq. 12, which is unacceptable and the details are in Appendix H. Thus, we propose our method to directly train the parameters of the global model on the server side.
@@ 3.5 Train the parameters of the global model
We use $\mathbb{E}\left[\mathbf{A}\right]$ denotes $\sum_{k}^{K}(\mathbf{A}_{k})$, $\mathbb{E}\left[\mathbf{B}\right]$ denotes $\sum_{k}^{K}(\mathbf{B}_{k})$, $\mathbb{E}\left[\mathbf{A}\otimes\mathbf{B}\right]$ denotes $\sum_{k}^{K}(\mathbf{A}_{k}\otimes\mathbf{B}_{k})$. Previous works [46, 49] approximate the expectation of Kronecker products by a Kronecker product of expectations $\mathbb{E}\left[\mathbf{A}\otimes\mathbf{B}\right]\approx\mathbb{E}\left[\mathbf{A}\right]\otimes\mathbb{E}\left[\mathbf{B}\right]$ with an assumption of $\mathbf{A}_{k_{l}}$ and $\mathbf{B}_{k_{l}}$ are independent, which is called Expectation Approximation (EA). However, it may lead to a biased global expectation. The details are discussed in Appendix I. Instead, we could construct a linear objective after aggregating the approximation of local posteriors via using block-diagonal empirical Fisher information matrices. We denotes $\bar{\mathbf{M}}$ as the matrix formula of $\bar{\bm{\mu}}=\mathrm{vec}(\bar{\mathbf{M}})$, and the optimal solution of $f(\bar{\bm{\mu}})$ is $\bar{\bm{\mu}}^{*}=\mathrm{vec}(\bar{\mathbf{M}}^{*})$. We construct $f(\bar{\bm{\mu}})$ as a multi-variates linear objective function. When $\bar{\bm{\mu}}=\bar{\bm{\mu}}^{*}$ is optimal solution, $f(\bar{\bm{\mu}})=\mathbf{o}$, where $\mathbf{o}$ is a vector with all zero. Note that
@@
The above equation is a quadratic objective, and it can be solved by modern optimization tools efficiently and conveniently.
Since the main objective of the above problem is both convex and Lipschitz smooth w.r.t $\mathrm{vec}(\bar{\mathbf{M}})$, we can use the gradient descent method to solve it with a linear convergence rate (See detailed proof in Appendix J). Here, we use automatic differentiation to calculate the gradient w.r.t. $\bar{\mathbf{M}}$.
@@ 3.6 Overall FedLPA algorithm and discussions
In summary, the proposed algorithm FedLPA follows the same paradigm as the standard one-shot federated learning framework. In FedLPA, the clients locally train their models to get $\mathbf{M}_{k}$ and calculate the local co-variance over its training dataset using the layer-wise Laplace approximation to compute $\mathbf{A}_{k},\mathbf{B}_{k}$, as shown in Algorithm 1 in Appendix A. Subsequently, each client transmits their local $\mathbf{A}_{k},\mathbf{B}_{k},\mathbf{M}_{k}$ to the server. Following Algorithm 2 in Appendix A, the server aggregates these contributions to obtain the global expectation, as described in Eq. 7, then trains the global model parameters, as outlined in Eq. 14. Thus, the transmitted data between the clients and the server is solely $\mathbf{A}_{k},\mathbf{B}_{k},\mathbf{M}_{k}$ without any extra auxiliary dataset and label information.
@@
Note that FedLPA can be directly adopted in most common scenarios. For the special case that the neural model has enormous single-layer weight parameters, how to extend our proposed FedLPA is discussed in Appendix K. The detailed privacy discussion of FedLPA could be found in Appendix L.
3.7 t-SNE observation and discussions
To quickly demonstrate the effectiveness of FedLPA, We show the t-SNE visualization of our FedLPA global model on the MNIST dataset as an example with a biased training data setting among 10 local clients. The experiment details, t-SNE visualizations of the local models and the global models of other algorithms and discussions are in Appendix M.2. As shown in Figure 1, FedLPA generates the global model which can clearly distinguish these classes, meanwhile, the classes are separate.
@@ 4.1 Experiments settings
Datasets. We conduct experiments on MNIST [55], Fashion-MNIST [56], CIFAR-10 [57], and SVHN [58] datasets [^1] In most of the previous works and the most popular benchmark, the majority of their experiments are using these datasets and these models. We choose these datasets and models to do the majority of our experiments following these established methods and benchmark to fairly compare our method with the baselines. The experiments with more complex datasets are in Appendix M.11.. We use the data partitioning methods for non-IID settings of the benchmark  [^2] https://github.com/Xtra-Computing/NIID-Bench to simulate different label skews. Specifically, we try two different kinds of partition: 1) #C = $k$: each client only has data from $k$ classes. We first assign $k$ random class IDs for each client. Next, we randomly and equally divide samples of each class to their assigned clients; 2) $p_{k}$ Dir($\beta$): for each class, we sample from Dirichlet distribution $p_{k}$ and distribute $p_{k,j}$ portion of class $k$ samples to client $j$. In this case, smaller $\beta$ denotes worse skews.
@@
Training Details. By default, we follow FedAvg [12] and other existing studies [59, 60, 15] to use a simple CNN with 5 layers in our experiments. The experiments with more complex neural network structures are in Appendix M.10. We set the batch size to 64, the learning rate to 0.001, and the $\lambda=0.001$ for FedLPA. By default, we set 10 clients and run 200 local epochs for each client. For the various settings of the number of clients and local epochs, we refer to Section 4.3 and Section 4.4. For results with error bars, we run three experiments with different random seeds. Note that all methods were evaluated under fair comparison settings. Due to the page limit, representative results are represented in the main paper. Refer to Appendix M for more experimental details and additional results.
@@
Baselines. To ensure fair comparisons, we neglect the comparison with methods that require to download auxiliary models or datasets, such as FedBE [61], FedKT [14] and FedGen [21], or even pretrained large model, like FedDISC [20]. FedOV [15] and FEDCAVE [23] entail sharing more client-side label information or transmitting client label information to the server, which could jeopardize label privacy and are beyond the scope of this study. XorMixFL [22] may not be practical, as we mentioned before. FedFisher [41] is not publicly available. FedDF [26], DAFL [35] and ADI [36] are compared with the state-of-the-art data-free method DENSE [24]. Co-Boosting [39] requires too many computational resources [^3] The experiments with more models, FedOV and Co-Boosting, are in Appendix M.7. In conclusion, we include one-shot FL algorithms as baselines including FedAvg [12], FedProx [31], FedNova [30], SCAFFOLD [29] and DENSE [24]. All the methods are fairly compared, and our implementation is available and the experiment details can be viewed in Appendix M.14.

### v5

@@ 1 Introduction
  • We are the first to train global model parameters via constructing a multi-variate linear objective function and optimizing its quadratic form, which allows us to formulate and solve the problem in a convex form efficiently, which has a linear convergence rate, ensuring good performance.
@@ 2.1 Federated learning on non-iid data
Previous work FedAvg [9] first introduced the concept of FL and presented the algorithm, which
achieved competitive performance on i.i.d data, in comparison to several centralized techniques. However, it was observed in previous works [27, 28] that the convergence rate and ultimate accuracy of FedAvg on non-IID data distributions were significantly reduced, compared to the results observed with homogeneous data distributions.
@@ 2.2 One-shot federated learning
One-shot Federated Learning (FL) is an emerging and promising research direction characterized by its minimal communication cost. In the first study on one-shot FL [13], the approach involved the aggregation of local models, forming an ensemble to construct the final global model. Subsequently, knowledge distillation using public data was applied in the following step. FedKT [14] brought forward the concept of consistent voting to fortify the ensemble. Recent research endeavors [19, 24] proposed data-free knowledge distillation schemes tailored for one-shot FL. These methods adopted the basic ensemble distillation framework as FedDF [26]. XorMixFL [22] introduced the use of exclusive OR operation (XOR) for encoding and decoding samples in data sharing. It is important to note that XorMixFL assumed the possession of labeled samples from a global class by all clients and the server, which might not align with practical real-world scenarios. A noteworthy innovation of DENSE [24] was its utilization of a generator to create synthetic datasets on the server side, circumventing the need for a public dataset in the distillation process. Co-Boosting [39] improves the ensemble when doing the distillation to improve the performance. FedOV [15] delved into addressing comprehensive label skew cases. FEDCVAE [23] confronted this challenge by transmitting all label distributions from clients to servers. These schemes [22, 14, 19, 24, 23, 15] exposed some client-side private information, leading to additional communication overhead and potential privacy leakage, e.g., FEDCVAE [23] needed all the client label distribution to be transmitted to the server side and FedOV [15] needed the clients to know the labels which were unknown. Instead, MA-Echo [40] adopted a unique approach by emphasizing the addition of norms among layer-wide parameters during the aggregation of local models. The project [41] focused on the theoretic analysis of the error in its approximation method. However, their method grappled with limited experiments and lacked detailed explanations of the approach. FedDISC [20], on the other hand, relied on the pre-trained model CLIP from OpenAI, where their reliance might not always align with practicality or suitability for diverse scenarios.
@@ 3.1 Objective formulation
Previous works [16, 17] give a common formula of the global posterior which consists of local posteriors $p(\bm{\theta}|\mathcal{D}_{k})$ under variational inference formulation.
@@ 3.2 Approximating posteriors
where $\bm{\theta}^{*}$ is the optimal parameter vector, $\bar{\mathbf{H}}=\mathbb{E}_{s\in\mathcal{D}}[\mathbf{H}]$ is the average Hessian of the negative log posterior over a dataset $\mathcal{D}$. It is reasonable to approximate global and local posteriors as multi-variates Gaussian distributions with expectations $\bar{\bm{\mu}}=\bm{\theta}^{*}$ and $\bm{\mu_{k}}=\bm{\theta}_{k}^{*}$; co-variances $\bar{\mathbf{\Sigma}}=\bar{\mathbf{H}}^{-1}$ and $\mathbf{\Sigma}_{k}=\bar{\mathbf{H}}_{k}^{-1}$ [44].
@@
As a result, if given local expectation $\bm{\mu}_{k}$ and local co-variance $\mathbf{\Sigma}_{k}$, the global posterior is determined by Eq. 2 as below:
@@
Therefore, the local co-variance $\mathbf{\Sigma}_{k}$ can be approximated by the empirical Fisher $\tilde{\mathbf{F}}_{k}$ [48, 49].
@@
The works [50, 51, 17] ignore co-relations between different parameters and only consider the self-relations of parameters as computing all co-relations is impossible. Thus, their methods are inaccurate. Detailed discussions and the novelty compared to previous works are in Appendix B.
In order to capture co-relations between different parameters efficiently, previous works [46, 43] estimate a block empirical Fisher information matrix $\mathbf{F}$ instead of assuming parameters are independent and approximating the co-variance by the diagonal of the empirical Fisher. As pointed out, co-relations inner a layer are much more significant than others [46, 52, 53], while computing the co-relations between different layers brings slight improvement but much more computation [54, 43]. Therefore, assuming parameters are layer-independent is a good trade-off. As a result, the approximated layer-wise empirical Fisher is block-diagonal. For layer $l$ on client $k$, its empirical Fisher $\mathbf{F}_{k_{l}}$ is one of the diagonal blocks in the whole empirical Fisher for the local model and is factored into two small matrices as below,
@@
where $\otimes$ is the Kronecker product; $\mathbf{A}_{k_{l}}=\mathbb{E}\left[\hat{\mathbf{a}}_{k_{l-1}}\hat{\mathbf{a}}_{k_{l-1}}^{\top}\right]+\pi_{l}\sqrt{\lambda}\mathbf{I}$ and $\mathbf{B}_{k_{l}}=\mathbb{E}\left[\hat{\mathbf{b}}_{k_{l}}\hat{\mathbf{b}}_{k_{l}}^{\top}\right]+\frac{1}{\pi_{l}}\sqrt{\lambda}\mathbf{I}$ are two expectation factor matrices over the data samples; $\hat{\mathbf{a}}_{k_{l}}$ is the activations and $\hat{\mathbf{b}}_{k_{l}}$ is the gradient of the pre-activations of layer $l$ on client $k$, $\lambda$ is the hyperparameter and $\pi_{l}$ is a factor minimizing approximation error in $\mathbf{F}_{k_{l}}$ [46, 49, 55]. $\mathbf{A}_{k_{l}}$ and $\mathbf{B}_{k_{l}}$ are symmetric positive definite matrices [45, 46].
@@ 3.4 Estimating the global expectation
As shown in Eq. 11, obtaining the global expectation $\bar{\bm{\mu}}_{l}$ requires calculating the inverse of $\bar{\mathbf{\Sigma}}_{l}^{-1}$ as Eq. 12, which is unacceptable and the details are in Appendix C. Thus, we propose our method to directly train the parameters of the global model on the server side.
@@ 3.5 Train the parameters of the global model
We use $\mathbb{E}\left[\mathbf{A}\right]$ denotes $\sum_{k}^{K}(\mathbf{A}_{k})$, $\mathbb{E}\left[\mathbf{B}\right]$ denotes $\sum_{k}^{K}(\mathbf{B}_{k})$, $\mathbb{E}\left[\mathbf{A}\otimes\mathbf{B}\right]$ denotes $\sum_{k}^{K}(\mathbf{A}_{k}\otimes\mathbf{B}_{k})$. Previous works [46, 49] approximate the expectation of Kronecker products by a Kronecker product of expectations $\mathbb{E}\left[\mathbf{A}\otimes\mathbf{B}\right]\approx\mathbb{E}\left[\mathbf{A}\right]\otimes\mathbb{E}\left[\mathbf{B}\right]$ with an assumption of $\mathbf{A}_{k_{l}}$ and $\mathbf{B}_{k_{l}}$ are independent, which is called Expectation Approximation (EA). However, it may lead to a biased global expectation. The details are discussed in Appendix D. Instead, we could construct a linear objective after aggregating the approximation of local posteriors via using block-diagonal empirical Fisher information matrices. We denotes $\bar{\mathbf{M}}$ as the matrix formula of $\bar{\bm{\mu}}=\mathrm{vec}(\bar{\mathbf{M}})$, and the optimal solution of $f(\bar{\bm{\mu}})$ is $\bar{\bm{\mu}}^{*}=\mathrm{vec}(\bar{\mathbf{M}}^{*})$. We construct $f(\bar{\bm{\mu}})$ as a multi-variates linear objective function. When $\bar{\bm{\mu}}=\bar{\bm{\mu}}^{*}$ is optimal solution, $f(\bar{\bm{\mu}})=\mathbf{o}$, where $\mathbf{o}$ is a vector with all zero. Note that
@@
The above equation is a quadratic objective, and it can be solved by modern optimization tools efficiently and conveniently. Since the main objective of the above problem is both convex and Lipschitz smooth w.r.t $\mathrm{vec}(\bar{\mathbf{M}})$, we can use the gradient descent method to solve it with a linear convergence rate. Here, we use automatic differentiation to calculate the gradient w.r.t. $\bar{\mathbf{M}}$.
Algorithm 1 FedLPA Global Aggregation
1:Input: clients $K$, layers $L$
2:Initialize global weight $\bar{\mathbf{W}}_{l}$ of layer $l=1,...,L$
3:clients executes:
4:Initialize local model
5:for k = 1, …, K do
7:end for
8:Server executes:
9:for l = 1, …, L do
10:     $\bar{\mathbf{A}}_{l}\leftarrow\sum_{k}^{K}\mathbf{A}_{k_{l}}$
11:     $\bar{\mathbf{B}}_{l}\leftarrow\sum_{k}^{K}\mathbf{B}_{k_{l}}$
14:     $\bar{\mathbf{W}}_{l}\leftarrow\bar{\mathbf{M}}_{l}$
15:end for
@@ 3.6 Overall FedLPA algorithm and discussions
In summary, the proposed algorithm FedLPA follows the same paradigm as the standard one-shot federated learning framework. In FedLPA, the clients locally train their models to get $\mathbf{M}_{k}$ and calculate the local co-variance over its training dataset using the layer-wise Laplace approximation to compute $\mathbf{A}_{k},\mathbf{B}_{k}$. Subsequently, each client transmits their local $\mathbf{A}_{k},\mathbf{B}_{k},\mathbf{M}_{k}$ to the server. Following Algorithm 1, the server aggregates these contributions to obtain the global expectation, as described in Eq. 7, then trains the global model parameters, as outlined in Eq. 14. Thus, the transmitted data between the clients and the server is solely $\mathbf{A}_{k},\mathbf{B}_{k},\mathbf{M}_{k}$ without any extra auxiliary dataset and label information.
Note that FedLPA can be directly adopted in most common scenarios. For the special case that the neural model has enormous single-layer weight parameters, how to extend our proposed FedLPA is discussed in Appendix E.
3.7 t-SNE observation and discussions
To quickly demonstrate the effectiveness of FedLPA, we show the t-SNE visualization of our FedLPA global model on the MNIST dataset as an example with a biased training data setting among 10 local clients. The experiment details, t-SNE visualizations of the local models and the global models of other algorithms and discussions are in Appendix G.1. As shown in Figure 1, FedLPA generates the global model which can clearly distinguish these classes, meanwhile, the classes are separate.
@@
3.8 Privacy Discussions
FedLPA is intuitively compatible with existing privacy-preserving techniques, such as differential privacy (DP) [56, 57], secure multiparty computation (SMC) [58, 59], and homomorphic encryption (HE) [60, 61, 62]. In Appendix F.1, we propose a naive DP-FedLPA with two different mechanisms to show the compatibility with differential privacy. Meanwhile, we mention that our proposed FedLPA has the same privacy-preserving level as the conventional federated learning algorithms (i.e, FedAvg, FedProx, FedNova and Dense). Compared with FedAvg, we have conducted a detailed analysis from a privacy attack perspective to show that our proposed FedLAP exhibits a security level consistent with FedAvg against several types of privacy attacks, where the details are shown in Appendix F.3. Note that the main focus of FedLPA is to improve the learning performance on the one-shot FL settings, thus, we leave the integration with other privacy-preserving techniques beyond DP as an open problem.
@@ 4.1 Experiments settings
Datasets. We conduct experiments on MNIST [63], Fashion-MNIST [64], CIFAR-10 [65], and SVHN [66] datasets. In most of the previous works and the most popular benchmark, the majority of their experiments use these datasets and these models. We choose these datasets and models to do the majority of our experiments following these established methods and benchmarks to fairly compare our method with the baselines. We use the data partitioning methods for non-IID settings of the benchmark  [^1] https://github.com/Xtra-Computing/NIID-Bench to simulate different label skews. Specifically, we try two different kinds of partition: 1) #C = $k$: each client only has data from $k$ classes. We first assign $k$ random class IDs for each client. Next, we randomly and equally divide samples of each class to their assigned clients; 2) $p_{k}$ Dir($\beta$): for each class, we sample from Dirichlet distribution $p_{k}$ and distribute $p_{k,j}$ portion of class $k$ samples to client $j$. In this case, smaller $\beta$ denotes worse skews.
Here’s a brief overview of these datasets. MNIST Dataset: The MNIST dataset comprises binary images of handwritten digits. It consists of 60,000 28x28 training images and 10,000 testing images. FMNIST Dataset: Similar to MNIST, the FMNIST dataset also contains 60,000 28x28 training images and 10,000 testing images. SVHN Dataset: The SVHN dataset includes 73,257 32x32 color training images and 10,000 testing images. CIFAR-10 Dataset: CIFAR-10 consists of 60,000 32x32 color images distributed across ten classes, with each class containing 6,000 images. The input dimensions for MNIST, FMNIST, SVHN, and CIFAR-10 are 784, 784, 3,072, and 3,072, respectively.
@@
Training Details. By default, we follow FedAvg [12] and other existing studies [67, 68, 15] to use a simple CNN with 5 layers in our experiments. The experiments with more complex neural network structures are in Appendix G.8. We set the batch size to 64, the learning rate to 0.001, and the $\lambda=0.001$ for FedLPA. By default, we set 10 clients and run 200 local epochs for each client. For the various settings of the number of clients and local epochs, we refer to Section 4.3 and Section 4.5. For results with error bars, we run three experiments with 5 different random seeds. Note that all methods were evaluated under fair comparison settings. Due to the page limit, representative results are represented in the main paper. Refer to Appendix G for more experimental details and additional results.
@@
Baselines. To ensure fair comparisons, we neglect the comparison with methods that require to download auxiliary models or datasets, such as FedBE [69], FedKT [14] and FedGen [21], or even pretrained large model, like FedDISC [20]. FedOV [15] and FEDCAVE [23] entail sharing more client-side label information or transmitting client label information to the server, which could jeopardize label privacy and are beyond the scope of this study. XorMixFL [22] may not be practical, as we mentioned before. FedFisher [41] is not publicly available. FedDF [26], DAFL [35] and ADI [36] are compared with the state-of-the-art data-free method DENSE [24]. Co-Boosting [39] requires too many computational resources [^2] The experiments with more models, FedOV and Co-Boosting, are in Appendix G.5.. In conclusion, we include one-shot FL algorithms as baselines including FedAvg [12], FedProx [31], FedNova [30], SCAFFOLD [29] and DENSE [24]. All the methods are fairly compared, and our implementation is available and the experiment details can be viewed in Appendix G.11.